# Visual Evaluative AI: A Hypothesis-Driven Tool with Concept-Based Explanations and Weight of Evidence

### v2

@@ 4.2 Results
| PCBM | ||||
| PCBM+WoE | ||||
| ResneXt50 | Original | |||
@@
| PCBM | ||||
| PCBM+WoE | ||||
| Resnet152 | Original | |||
@@
Table 1 reports the performance of ICE(8) and ICE(8)+WoE using three different CNN backbone models (Resnet50, Resnet152 [6] and ResneXt50 [16]). ICE(8)+WoE and PCBM+WoE achieve comparable performance compared to the original CNN model. Moreover, ICE(8)+WoE and PCBM+WoE have a smaller standard deviation than the original ICE(8) and PCBM models, respectively. Since ICE uses the weights of the original CNN backbone models, we conduct a further ablation experiment by replacing the classifier layer of ICE with the Gaussian Naive Bayes (ICE+GNB) to compare with ICE+WoE. The results in Table 1 show that ICE(8)+WoE and ICE(8)+GNB have closer performance than ICE(8)+WoE and the original ICE(8). The reason is that our implementation using WoE and GNB are both Naive Bayes methods so they use similar loss functions when learning the concept scores.
Having more concepts does not lead to better accuracy
Figure 4 shows the performance of the original ResneXt50, ICE and ICE+WoE over different number of concepts from 5 concepts to 40 concepts. Two figures from the left show the performance of ICE using NMF reducer. When there are 5 concepts, ICE(5)+WoE () has a significantly higher F1-score than ICE(5) () (, ). Since we have 2048 features at the classifier layer of ResneXt50, ResneXt50 outperforms ICE(5)+WoE and ICE(5) significantly (). But the performance of both ICE+WoE and ICE match the performance of the original ResneXt50 when we have at least 7 concepts. Particularly, with as few as 7 concepts, ICE and ICE+WoE achieve similar performance to the original ResneXt50 using 2048 features. The performance of ICE and ICE+WoE also stopped improving at 7 concepts with a backbone of ResneXt50. The reason is that when we apply a reducer in ICE (e.g. NMF), some important concepts are detected at first. Then after we increase the number of concepts, some noisy concepts are detected, which could lead to a slight drop in the performance. Eventually, all important concepts are found and match the performance of the original CNN model. Similar to the findings in [19], when we compare the performance using different reducers, NMF and PCA (principal component analysis), PCA provided the best performance but could be less interpretable compared to NMF.
In summary, the results show that with a few number of concepts (e.g., 8 concepts in the demo Figure 3), we can achieve comparable performance compared to the original models. Therefore, this indicates the accuracy of the evidence being generated, which is potentially useful to the decision-makers. Moreover, it would also be easier for users to evaluate the evidence by not showing too many concepts.
@@ 5 Conclusion
In this paper, we introduce Visual Evaluative AI 333https://github.com/thaole25/VisualEvaluativeAI, a tool for hypothesis-driven decision support. This tool can highlight the high-level concepts in an image and provide positive/negative evidence for all possible hypotheses. Our tool is further applied and evaluated in the skin cancer domain with a web-based application called EvaSKan that offers skin cancer diagnosis support. In future work, a more comprehensive evaluation will be undertaken by addressing the domain expert opinions on this application.

### v3

@@ 4.2 Results
| PCBM(12) | ||||
| PCBM(12)+WoE | ||||
| ResneXt50 | Backbone | |||
@@
| PCBM(12) | ||||
| PCBM(12)+WoE | ||||
| Resnet152 | Backbone | |||
@@
Table 1 reports the performance of ICE(7), ICE(7)+WoE, PCBM(12) and PCBM(12)+WoE using three different CNN backbone models (Resnet50, Resnet152 [6] and ResneXt50 [16]). We select 12 concepts for PCBM based on previous work [18, 17]. For ICE, we run experiments with the number of concepts ranging from 5 to 40. As shown in Figure 4, performance peaks at 7 concepts. Therefore, the final comparison in this table is made between ICE(7) and PCBM(12).
The results show that ICE(7)+WoE and PCBM(12)+WoE achieve comparable performance to the original CNN models. Particularly, with ResneXt50, the F1-score of ICE(7)+WoE and PCBM(12)+WoE are and , respectively, while the original ResneXt50 has an F1-score of . Therefore, ICE(7)+WoE (using 7 features) and PCBM(12)+WoE (using 12 features) show comparable performance compared to the original ResneXt50 with 2048 features. Moreover, similar to the findings in [19], when we compare the performance using different reducers as in Figure 4, NMF and PCA (principal component analysis), PCA provided the best performance but could be less interpretable compared to NMF.
Having more concepts did not lead to better accuracy
Figure 4 shows the performance of the original ResneXt50, ICE and ICE+WoE over different numbers of concepts from 5 concepts to 40 concepts. Two figures from the left show the performance of ICE using the NMF reducer. When there are 5 concepts, ICE(5)+WoE () has a significantly higher F1-score than ICE(5) () (, ). Since we have 2048 features at the classifier layer of ResneXt50, ResneXt50 outperforms ICE(5)+WoE and ICE(5) significantly (). But the performance of both ICE+WoE and ICE match the performance of the original ResneXt50 when we have at least 7 concepts. Particularly, with as few as 7 concepts, ICE and ICE+WoE achieve similar performance to the original ResneXt50 using 2048 features. The performance of ICE and ICE+WoE also stopped improving at 7 concepts with a backbone of ResneXt50. The reason is that when we apply a reducer in ICE (e.g. NMF), some important concepts are detected at first. Then after we increase the number of concepts, some noisy concepts are detected, which could lead to a slight drop in the performance. Eventually, all important concepts are found and match the performance of the original CNN model.
In summary, the results show that with a few number of concepts (i.e., 7 concepts), we can achieve comparable performance compared to the original CNN models. Therefore, this indicates the accuracy of the evidence being generated, which is potentially useful to the decision-makers. Importantly, despite the concept-based models (ICE(7), ICE(7)+WoE, PCBM(12) and PCBM(12)+WoE) being slightly less accurate than the CNN backbones, it would also be much easier for users to interpret and evaluate the evidence by not showing too many concepts.
@@ 5 Conclusion
In this paper, we introduce Visual Evaluative AI 333https://github.com/thaole25/EvaluativeAI, a tool for hypothesis-driven decision support. This tool can highlight the high-level concepts in an image and provide positive/negative evidence for all possible hypotheses. Our tool is further applied and evaluated in the skin cancer domain with a web-based application called EvaSKan that offers skin cancer diagnosis support. In future work, a more comprehensive evaluation will be undertaken by addressing the domain expert opinions on this application.